differentiate with respect to x
x^(sin x – cos x) + (x)^2 – 1/ (x)^2 + 1

Grade:10

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ \\\frac{d}{dx}\left(x^{\left(\sin \left(x\right)-\cos \left(x\right)\right)}+x^2-\frac{1}{x^2}+1\right) \\=2x+\frac{2}{x^3}+x^{\sin \left(x\right)-\cos \left(x\right)-1}\left(\sin \left(x\right)\left(x\ln \left(x\right)+1\right)+\cos \left(x\right)\left(x\ln \left(x\right)-1\right)\right) \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty